OpenStudy (anonymous):

What is the difference of x plus 2 all over the quantity x minus 4, quantity squared. − x over x minus 4.?

OpenStudy (anonymous):

\[\frac{ x+2 }{ (x-4)^{2} }-\frac{ x }{x-4}\]

OpenStudy (anonymous):

Do you want to combine like terms or what?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well I think @Dtan1293

OpenStudy (luigi0210):

Easy, get common denominators then continue.

OpenStudy (anonymous):

so then multiply (x-4) by the numerator and denominator on the right hand side

OpenStudy (anonymous):

just on the right hand side?

OpenStudy (anonymous):

im lost when it comes to this

OpenStudy (luigi0210):

@Dtan1293 you can't do that. It's an Expression not an equation

OpenStudy (luigi0210):

You have to get common denominators

OpenStudy (anonymous):

okay so how do I do that?

OpenStudy (luigi0210):

Multiply the second by another x-4, top and bottom

OpenStudy (anonymous):

okay ss it is now \[\frac{ x+2 }{ (--4)^{2} }-\frac{ x-4(x) }{ x-4(x-4) }\] ?

OpenStudy (anonymous):

so

OpenStudy (anonymous):

(x-4)^2

OpenStudy (luigi0210):

There you go

OpenStudy (anonymous):

then what?

OpenStudy (luigi0210):

Distribute the top then do the needed operation

OpenStudy (anonymous):

okay thankyou